OpenStudy (anonymous):

find and classify the critical point of f(t)=t(t^2-4)^1/3 and find extreme values. i need the solution urgent.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

\[f(t)=t(t ^{2}-4)^{\frac{ 1 }{ 3 }}\] find critical point. i need its solution urgent help me!

OpenStudy (anonymous):

find f'(t),using D(uv)=uv'+u'v then equate it to 0 to find the points then find extreme values at these points.

OpenStudy (anonymous):

what is the answer??? \[\sqrt{12} or \pm \sqrt{\frac{ 6 }{ 5 }}\] ???? ans me quickly